OpenStudy (anonymous):

Find the simple interest and the total amount due on the loan. Principal Rate Time $22,400 5 1/2% 9 months

OpenStudy (ash2326):

Simple Interest=P*R*T/100 =22,400*11/2*9/12/100 t=9/12 year =22400*5.5*0.75/100 Simple Interest=$924 Total Amount=Principal+Simple Interest=$23324